OpenStudy (deadshot):

What is the simplified form of the quantity x-squared minus 16 over the quantity x plus 4 ?

OpenStudy (deadshot):

\[\frac{ x^2-16 }{ (x+4) }\]

geerky42 (geerky42):

\[\dfrac{x^2 - 16}{x+4}\] I'd factor the numerator if I were you.

geerky42 (geerky42):

Remember the differences of squares? \[x^2 - a^2 = (x-a)(x+a)\]

OpenStudy (deadshot):

so then, the numerator would be \[(x+4)(x-4)\]right?

geerky42 (geerky42):

Right.

OpenStudy (deadshot):

and then the (x+4) would cancel each other out

geerky42 (geerky42):

Yes

OpenStudy (deadshot):

leaving the answer as x - 4, with the restriction x ≠ -4

geerky42 (geerky42):

Yeah.

OpenStudy (deadshot):

Thanks!

geerky42 (geerky42):

Glad I helped.

OpenStudy (anonymous):

x-4, I did that one in my head

geerky42 (geerky42):

Good for you.